Entries for Issuing Bonds and Amortizing Discount by Straight-Line Method

On the first day of its fiscal year, Chin Company issued $14,300,000 of five-year, 7% bonds to finance its operations of producing and selling home improvement products. Interest is payable semiannually. The bonds were issued at a market (effective) interest rate of 9%, resulting in Chin Company receiving cash of $13,168,515.

a. Journalize the entries to record the following:

  1. Issuance of the bonds.
  2. First semiannual interest payment. The bond discount amortization, using the straight-line method, is combined with the semiannual interest payment. (Round your answer to the nearest dollar.)
  3. Second semiannual interest payment. The bond discount amortization, using the straight-line method, is combined with the semiannual interest payment. (Round your answer to the nearest dollar.)

For a compound transaction, if an amount box does not require an entry, leave it blank. Round your answers to the nearest dollar.

1.
2.
3.

b. Determine the amount of the bond interest expense for the first year.
$

c. Why was the company able to issue the bonds for only $13,168,515 rather than for the face amount of $14,300,000?
The market rate of interest is the contract rate of interest.
Answer #1
a
1 Cash 13168515
Discount on Bonds payable 1131485
       Bonds payable 14300000
2 Interest expense 613649
       Discount on Bonds payable 113149 =1131485/5*6/12
       Cash 500500 =14300000*7%*6/12
3 Interest expense 613649
       Discount on Bonds payable 113149
       Cash 500500
b
Bond interest expense for the first year 1227298 =613649+613649
c
The market rate of interest is greater than the contract rate of interest